OpenStudy (coconut):

1. What is the 19th term of the arithmetic sequence 15, 23, 31, ...? 2. What is the 25th term of the arithmetic series 9, 15, 21, ...?

OpenStudy (anonymous):

do you know the formula?

OpenStudy (coconut):

no sorry

OpenStudy (anonymous):

we start with 15. then for each subsequent term we add 8. so, for the 2nd term we add one 8 and for the third term we add two 8's to the starting point 15. so for the 19th term we will add 18 8's or 18*8=144 to 15 or 159. now can you do #2?

OpenStudy (anonymous):

\[\LARGE a_n=a_1+(n-1)d\] a_ is the first term and d is difference we have: \[\LARGE d=a_n-a_{n-1}\] \[\LARGE d=a_2-a_1\] \[\LARGE d=23-15\] \[\LARGE d=8\] now we substitute: \[\LARGE a_{19}=15+(19-1)8\] go on...

OpenStudy (coconut):

for number one i got 159 and number 2 i got 153

OpenStudy (anonymous):

those are correct